Osborn Manufacturing uses a predetermined overhead rate of $19.00 per direct labor-hour. This predetermined rate was based on a cost formula that estimates $243,200 of total manufacturing overhead for an estimated activity level of 12,800 direct labor-hours.


The company actually incurred $241,000 of manufacturing overhead and 12,300 direct labor-hours during the period.

Required:

1. Determine the amount of underapplied or overapplied manufacturing overhead for the period.

2. Assume that the company's underapplied or overapplied overhead is closed to Cost of Goods Sold. Would the journal entry to dispose of the underapplied or overapplied overhead increase or decrease the company’s gross margin? By how much?

by 1. Manufacturing overhead 2. The gross margin would by
Answer #1

Applied overhead = 12300*19 = 233700

Actual overhead = 241000

1) Manufacturing overhead under applied by (241000-233700) = 7300

If Under applied overhead closed in cost of goods sold then cost of goods sold increase and when cost of goods sold increase gross profit will be decrease

2) The gross margin would decrease by $7300